find 4 solutions for 7x+2y=5
Answer by elima(1433)

For this equation you make a table choosing values for x and solving the equation for y;
x|y
---
-1|
3|
-3|
So when x=-1;
7(-1)+2y=5
-7+2y=5
2y=5+7
2y=12
y=6
X=-1, y=6
7(3)+2y=5
21+2y=5
2y=5-21
2y=-16
y=-8
So when x=3, y=-8
7(-3)+2y=5
-21+2y=5
2y=5+21
2y=26
y=13
When x=-3, y=13
Hope you understand
=)